A goldsmith has two alloys that are different purities of gold. The first is three-fourths pure gold and the second is five twelfths pure gold. How many ounces of each should be melted and mixed in order to obtain a 6-oz mixture that is two-thirds pure gold?
Answer by checkley77(12844)

.75x+5/12(6-x)=2/3*6
.75x+5*6/12-5/12x=12/3
.75x+30/12-5x/12=4
.75x-5x/12+2.5=4
(.75x*12-5x)/12=4-2.5
(9x-5x)/12=1.5
4x=1.5*12
4x=18
x=18/4
x=4.5 oz. of 3/4 pure gold is used.
6-4.5=1.5 oz. of 5/12 pure gold is used.
Proof:
.75*4.5+5/12*1.5=2/3*6
3.375+.625=4
4=4